calculate the percent (m/m) concentration off a solution that contains 11 g of pure sodium hydroxide in 75g of solution
1
Expert's answer
2020-11-11T10:36:54-0500

11g/ 75g = x/ 100g

x = 100× 11 / 75

x = 14.67g


So concentration = 14.67g per 100g = 14.67 % (m/m)